OpenStudy (anonymous):

Find the center of the circle with equation x^2+y^2-4x+6y+1=0

OpenStudy (callisto):

For equation of circle in this form: x^2 + y^2 +Dx +Ey +F =0 centre = (-D/2 , -E/2) Now, D= -4, E = 6, can you find the answer?

OpenStudy (anonymous):

-2,-3?

OpenStudy (callisto):

Nope, y-coordinate is correct, not for the x.

OpenStudy (callisto):

for x coordinate, it's equal to -(-4)/2 =4/2 =?

OpenStudy (anonymous):

Ok! Makes sense now. Thx!